Please help me wnswer this problem. Sam's age 9years from now is three times her age 15 years ago
Answer by FrankM(1040)

S+9=3(S-15)
S+9=3S-45
54=2S
S=27
In 9 years S is 36 which is 3 times age 15 years ago, 12